OpenStudy (anonymous):

14x+18y= -54 -4x-14y=42

OpenStudy (savvy):

multiply the first eqn. by 2 and 2nd by 7....eliminate x and get y....now put the value of y you got in any of the eqn. and get x...

OpenStudy (anonymous):

is this method substitution?

OpenStudy (callisto):

That's elimination

OpenStudy (anonymous):

I need the process in substitution.

OpenStudy (callisto):

Then consider equation (2) -4x-14y=42 , make x as subject. -4x-14y=42 -4x = 42 + 14y x = (42 + 14y) / -4 -(3) Sub (3) into the first equation. Can you do it from here?

OpenStudy (savvy):

find out y in terms of x from the first eqn. then put it in the second one......get x....put the value of x u gt in any of the eqn. and get y...

OpenStudy (anonymous):

Hi Cllisto, you wrote the -4x-14y=42 twice, is that correct ?

OpenStudy (callisto):

lol, The first one is to tell you what to do... ''-4x-14y=42<- make x as subject. '' Then the rest are the steps

OpenStudy (anonymous):

lol ok